7. (12pts) Solve the equations a. 2sino - 1 = 0 b. 3cscd = 1
Answer #1

7) a) 2 Sino – 1 = 0 2 Sino - Sino = not- 2 T = Sin T Sino TE TT Ⓡ = nte + Te+ 6 = 0, +1, +2,- Sino Sina = +61996 Whire mz0,